OpenStudy (anonymous):

Craig bought a second hand-hand bicycle for $80 and after repairing it sold it for $110. What is his profit as a percentage of the cost?

OpenStudy (anonymous):

profit=(110-80)/80=37.5%

OpenStudy (anonymous):

yes, the profit is 110-80 which is 30. Now what percentage of 80 is 30? That gives 30/80 * 100 = 37.5%